OpenStudy (anonymous):

P(t) = 6e^0.2t/(4+2e^0.2t) determine average rate of change of the population for 0<=t<=20 Find a time at which instantaneous rate of change of population is equal to the average rate of change

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@dumbcow

OpenStudy (swissgirl):

Firstly find f(0) and f(20) Then we need to find the average rate of change from 0<t<20 So we would use the following formula \[\text {Average Rate of Change } =\frac{f(20)-f(0)}{20-0}\]

OpenStudy (anonymous):

Ok yes and I got like .0947

OpenStudy (dumbcow):

set that equal to P'(t) ..... solve for t

OpenStudy (dumbcow):

use quotient rule for finding derivative

OpenStudy (fibonaccichick666):

You can always relate rate of change to derivative(side note). It essentially implies take the derivative :)

OpenStudy (anonymous):

So check is this right 1.2e^.2t/(4+2e^.2)^2

OpenStudy (anonymous):

Opps Should be 1.2e^.2t/(4+2e^.2t)^2

OpenStudy (anonymous):

Is that right ?

OpenStudy (dumbcow):

pretty close ... a 2 can be factored out of denominator

OpenStudy (dumbcow):

1.2e^.2t/(2+e^.2t)^2

OpenStudy (anonymous):

So now how do I solve when set equal to .0947 ? Cross multiply ?

OpenStudy (dumbcow):

yeah, looks like you will need to distribute it out and get a quadratic

OpenStudy (anonymous):

So i get 0.3788e^.2t+.0947e^.4t+0.3788=1.2e^.2t

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

hm?

OpenStudy (anonymous):

How do now solve this

OpenStudy (fibonaccichick666):

for instantaneous? get the derivative and evaluate the difference between 0 and 20

OpenStudy (fibonaccichick666):

sorry misread, no just do what dumbcow did

OpenStudy (anonymous):

But thats where I went but how do I solve from there ?

OpenStudy (fibonaccichick666):

combine the .2 terms to get a quadratic then use e^.4t=x^2

OpenStudy (fibonaccichick666):

That I think is what I'd do

OpenStudy (anonymous):

Ah :)

OpenStudy (anonymous):

?? I am getting imaginary numbers

OpenStudy (anonymous):

I did ln and got a number thanks

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Spectrum

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

Please help I just want to finish this one

OpenStudy (anonymous):

other people already told you how to do it. Just find the slope

OpenStudy (anonymous):

So I found derivitave an set it equal to the 0.0947 how now do I solve for t

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

How do I solve this help!

OpenStudy (anonymous):

:(